# Solving stochastic climate-economy models: A deep least-squares Monte Carlo approach

### v1

@@ 2.1 Modeling uncertainty
Stochastic process uncertainty refers to the uncertainty in the evolution of future trajectories of exogenous variables. A classical example from quantitative finance is Brownian motion, $B=(B_{t})_{t\geq 0}$, modeled by $B_{0}=0$ and $B_{t+h}-B_{t}\sim\mathcal{N}(0,h)$ for $t\geq 0$ and $h\geq 0$, where $\mathcal{N}(0,h)$ denotes the normal distribution with expected value 00 and variance $h$. Incorporating stochastic process uncertainties is challenging because the uncertainty propagates over time, increasing the volatility of the variable’s distribution. The LSMC method we present below is highly sensitive to introduced volatility, making this incorporation a significant challenge that few contributions in the climate-economy literature have successfully addressed.
@@
Initial parameter uncertainty refers to uncertainty about one or more parameters in the system that remain fixed over time. A common method to study this uncertainty is a perturbation analysis, where parameters are sampled, the model is solved, and the process is repeated. However, this approach does not accurately depict the model’s evolution over time, as an agent in the model would consider overall outcome uncertainty, not individual instances of the uncertain parameter. Another related concept is Bayesian learning (Kelly and Kolstad, 1999), where the parameter distribution evolves over time as more information about the system is revealed. This type of uncertainty can be treated by the LSMC approach presented in this paper, but we chose not to include this in the current study, leaving it for future work.
@@
Identifying reasonable uncertainties to include in the model is challenging, as some uncertainties might be more significant than others. Advanced statistical analyses are required to make educated assumptions about probability distributions for the climate and economc system. For our paper, we incorporate five uncertainties into the DICE model, as identified by Nordhaus (2018). These include stochastic process uncertainties in the growth rates of total factor productivity $A$ and the rate of decarbonization $\sigma$, as well as initial parameter uncertainties in the temperature-sensitivity coefficient, the damage coefficient, and the carbon cycle coefficient. We emphasize that our method is not limited to these specific uncertainties, and we now explain our choices in more detail.
@@
Productivity growth. Assuming a Cobb-Douglas production function, the growth in total factor productivity $A$ models the growth in output that is not explained by growth in inputs of labor and capital used in production. The DICE model assumes $A$ evolves according to $A_{t+1}=A_{t}/(1-g_{A}(t))$, where $g_{A}(t)$ is the deterministic growth rate which is specified in Table 1. Nordhaus (2018) assumes $g_{A}(0)$ is normally distributed with mean 0.07600760.0760.076 and standard deviation 0.05600560.0560.056. But in this case, using the dynamics for the growth rate, we can model $g_{A}(t)$ as normally distributed with mean $g_{A}(0)\exp(-0.005t\Delta)$ and standard deviation $0.056\exp(-0.005t\Delta)$. In order to remove extreme cases, we truncate this distribution at the mean $\pm$ two standard deviations. The evolution of $A_{t}$ is shown in Figure 1.
@@
Abbildung 1: Evolution of total factor productivity $A$ under the assumption that the growth rate $g_{A}$ is uncertain.
@@
The rate of decarbonization. Uncontrolled industrial $\mathrm{CO}_{2}$ emissions are given by a level of carbon intensity, $\sigma_{t}$, times gross output. The DICE model assumes $\sigma$ evolves according to $\sigma_{t+1}=\sigma_{t}\exp(g_{\sigma}(t)\Delta)$, with a deterministic growth rate $g_{\sigma}(t)$ which is specified in Table 1. Nordhaus (2018) assumes $g_{\sigma}(0)$ is normally distributed with mean $-0.0152$ and standard deviation 0.0032000320.00320.0032. We therefore model $g_{\sigma}(t)$ as normally distributed with mean $g_{\sigma}(0)(1-0.001)^{t\Delta}$ and standard deviation $0.0032(1-0.001)^{t\Delta}$, truncating the distribution at the mean $\pm$ two standard deviations in order to remove extreme cases. The evolution of $\sigma_{t}$ is shown in Figure 2.
@@
Equilibrium temperature sensitivity (ETS). The equilibrium temperature sensitivity measures how much the Earth’s surface will warm in response to a doubling of atmospheric $\mathrm{CO_{2}}$. The DICE model assumes the ETS is equal to $3.1\degree\mathrm{C}$ for an equilibrium $\mathrm{CO_{2}}$ doubling. In Table 1, the ETS corresponds to the denominator in the definition of $\psi_{2}$. Nordhaus (2018) models the ETS as a log-normal distribution, $\exp(X)$ with $X\sim\mathcal{N}(1.1060,0.2646^{2})$. We do the same, truncating at the mean $\pm$ two standard deviations.
The damage function. The DICE model assumes climate-induced economic damages are a quadratic function of the increase in atmospheric temperature. It is modeled as a fractional loss of global output from greenhouse warming, $d(t)=\pi_{2}\times(T_{t}^{\mathrm{AT}})^{2}$, where $\pi_{2}$ denotes a damage coefficient representing the severity of the economic impact of global warming. The DICE model assumes $\pi_{2}$ to be equal to 0.00236. Nordhaus (2018) models the $\pi_{2}$ by a normal distribution with mean 0.00236 and standard deviation 0.00118. We use the same distribution but truncate it at the mean minus one standard deviation, and at the mean plus two standard deviations.
The carbon cycle. The carbon cycle coefficient models the equilibrium concentration of carbon in the biosphere and upper level of the oceans. The DICE model assumes it to be equal to 360 gigatonnes of carbon (GtC). In Table 1, it corresponds to the value 360 appearing in the definitions of $\phi_{12}$ and $\phi_{23}$. Nordhaus (2018) models this coefficient as a log-normal distribution, $\exp(X)$ with $X\sim\mathcal{N}(5.8510,0.2649^{2})$ We do the same, truncating at the mean $\pm$ two standard deviations.
Abbildung 3: Density plots of the parameter distributions of equilibrium temperature sensitivity (left panel), the damage coefficient (middle panel), and carbon cycle coefficient (right panel).
@@
  • Assuming $\pi_{2}\sim\mathcal{N}(0.00236,0.00118^{2})$ implies a roughly $2.3\%$ probability of $\pi_{2}$ being negative. This is a non-negligible scenario. Given that the DICE model aims to combine equations for the economy and climate, it is highly questionable to assume the damage coefficient could be below or just above zero. Moreover, the assumption of a log-normal distribution for the equilibrium temperature sensitivity and the carbon cycle coefficient also entails a non-negligible probability of those coefficients being close to zero.
    Nordhaus (2018) avoids this issue by discretizing the distributions, separating them into quintiles, and then calculating the expected values of the random variables within those quintiles. These expected values are taken as realizations of discrete uncertain variables, yielding sufficiently positive lowest realizations for the coefficients. Inspired by this approach, we also truncate the distributions of the random variables, however, without discretizing them. This avoids issues with too low damage coefficients and temperature sensitivities, as well as extreme growth rates for total factor productivity and carbon intensity.
@@ 3 The deep least-squares Monte Carlo method
The numerical solution of the model is achieved using the endogenous state least-squares Monte Carlo (LSMC) algorithm with control randomization, as introduced by Kharroubi et al. (2014) and adapted for expected utility optimal stochastic control problems by Andréasson and Shevchenko (2022). This method approximates the conditional expectation of the value function in the Bellman equation using regression with a quadratic loss function applied to the transformed value function. Typically, regression basis functions are ordinary polynomials of the state and control variables, usually up to the third order. In our implementation, we use deep neural networks to approximate the regression predictor. To mitigate transformation bias in the regression estimate of the conditional expectation, we employ the smearing estimate as proposed by Andréasson and Shevchenko (2022). Below is a brief description of the LSMC algorithm.
@@
where $Z_{0},Z_{1},\ldots,Z_{N-1}$ are independent disturbance terms, i.e. the state of the next period depends on the current state’s value, the current period’s control decision, and the realisation of the disturbance term.
@@
[…]
where the expectation is conditional on the state $X_{t}=x$ and the policy $\pi_{t}$ at time $t$. For further details on dynamic programming, we refer the interested reader to the excellent monograph by Fleming and Soner (2006) on the subject.
@@
Here, $\mathcal{A}_{t}$ denotes a set of admissible values of $\pi_{t}$, which may depend on $x$. When the number of state variables is more than three, it usually becomes computationally infeasible to use quadrature-based methods to evaluate the conditional expectation in (8), making simulation methods like LSMC preferable.
@@ 3.1 Transformation bias and heteroskedasticity
In Equation (16), the mean of the transformed residuals does not depend on $(X_{t},\pi_{t})$, simplifying the function evaluation of $\hat{\Phi}_{t}(X_{t},\pi_{t})$, as the mean can be precomputed and reused.
If heteroskedasticity is present in the regression with respect to the state and control variables, a method that accounts for heteroskedasticity is required. In this case, the conditional variance can be modelled as a function of covariates:
@@
where $\{\bm{g}_{\theta}\colon\theta\in\widehat{\Theta}_{t}\}$ is another parametrized family of predictor functions. There are various standard methods to estimate $\bm{g}_{\theta}$ and the smearing estimate with controlled heteroskedasticity can then be used as discussed in Andréasson and Shevchenko (2022).
@@
  • The method presented in Algorithm 1 is called the regression surface approach. A common alternative is the realized value approach, where the value function $V_{t}(x)$ in Equation (8) is not computed by using the approximation of the conditional expectation (which was needed to find the optimal policy according to Equation (9)), but rather by computing the discounted sum of rewards along one trajectory starting from the state $x$ at time $t$. While promising greater numerical stability than the regression surface approach, the realized value approach requires calculating optimal decisions along the individual trajectories, which comes at a significant computational cost. For details on this approach, we refer to Andréasson and Shevchenko (2022) and references therein. Originally, we also implemented the realized value approach, however, we found that the regression surface approach provided a sufficiently accurate solution for the number of sample points chosen in our numerical study in Section 4.
@@ 3.2 Neural networks
In our paper, we choose for the parametrized family of functions $\{\bm{f}_{\theta}\colon\theta\in\Theta\}$ the class of deep neural networks. This algorithmically generated class of functions has found tremendous success in all fields of science. Over the years, it has been shown that neural networks can act as surrogate functions in many models, due to their far reaching approximation capabilities.
Theorems that establish approximations are referred to as universal approximation theorems (UAT); notable contributions include Cybenko (1989) and Hornik (1991). These theorems establish the topological density of sets of neural networks in various topological spaces. One speaks of the universal approximation property (Kratsios, 2021) of a class of neural networks. Unfortunately, these theorems are usually non-constructive. To numerically find optimal neural networks, one typically combines backpropagation (see, for example, Rumelhart et al. (1986)) with ideas from stochastic approximation (Robbins and Monro, 1951; Kiefer and Wolfowitz, 1952; Dvoretzky, 1956).
Assuming sufficient integrability, the conditional expectation in Equation (10) is the orthogonal projection of $\rho V_{t+1}(X_{t+1})$ onto the subspace spanned by $(X_{t},\pi_{t})$ in the space of square-integrable random variables. The universal approximation property of neural networks in this space (see, for instance, Hornik (1991, Theorem 1)) then justifies the approximation of $\Phi_{t}(X_{t},\pi_{t})$ by $\bm{f}_{\theta}(X_{t},\pi_{t})$ for a suitably chosen neural network $\bm{f}_{\theta}$.
@@
6:         sample $Z_{t}^{m}$ from the distribution specified by the $\triangleright$ Disturbance
@@
4:         $\widehat{V}_{t}(\widetilde{X}_{t})\coloneqq R_{N}(\widetilde{X}_{t})$

### v2

@@ 2.1 Modeling uncertainty
Stochastic process uncertainty refers to the uncertainty in the evolution of future trajectories of exogenous variables. A classical example from quantitative finance is Brownian motion, $B=(B_{t})_{t\geq 0}$, modeled by $B_{0}=0$ and $B_{t+h}-B_{t}\sim\mathcal{N}(0,h)$ for $t\geq 0$ and $h\geq 0$, where $\mathcal{N}(0,h)$ denotes the normal distribution with expected value 0 and variance $h$. Incorporating stochastic process uncertainties is challenging because the uncertainty propagates over time, increasing the volatility of the variable’s distribution. The LSMC method we present below is highly sensitive to introduced volatility, making this incorporation a significant challenge that few contributions in the climate-economy literature have successfully addressed.
@@
Initial parameter uncertainty refers to uncertainty about one or more parameters in the system that remain fixed over time. A common method to study this uncertainty is a perturbation analysis, where parameters are sampled, the model is solved, and the process is repeated. However, this approach does not accurately depict the model’s evolution over time, as an agent in the model would consider overall outcome uncertainty, not individual instances of the uncertain parameter. Another related concept is Bayesian learning (kelly1999), where the parameter distribution evolves over time as more information about the system is revealed. This type of uncertainty can be treated by the LSMC approach presented in this paper, but we chose not to include this in the current study, leaving it for future work.
@@
Identifying reasonable uncertainties to include in the model is challenging, as some uncertainties might be more significant than others. Advanced statistical analyses are required to make educated assumptions about probability distributions for the climate and economic system. For our paper, we incorporate five uncertainties into the DICE model, as identified by nordhaus2018. These include stochastic process uncertainties in the growth rates of total factor productivity $A$ and the rate of decarbonization $\sigma$, as well as initial parameter uncertainties in the temperature-sensitivity coefficient, the damage coefficient, and the carbon cycle coefficient. Other uncertainty dimensions studied in the IAM literature include structural/model uncertainty and ambiguity / robust-control formulations; see, e.g., barnett2020pricing. We emphasize that our method is not limited to the chosen uncertainties, and we now explain our choices in more detail. Unless stated otherwise, we assume that the two process shocks driving productivity and decarbonization are independent across time and mutually independent, and that they are independent of the three time-0 parameter draws. The parameter draws are realized at $t=0$ and then held fixed along each simulated trajectory.
Remark 2.4.
We adopt the parametric uncertainty inputs proposed by nordhaus2018 primarily as a transparent benchmark that facilitates comparison with the existing stochastic-DICE literature, and not as a claim that these inputs represent a definitive or up-to-date probabilistic calibration. Recent IPCC assessments summarize physical-science evidence largely in terms of assessed ranges and scenario ensembles (SSP/RCP), rather than as a ready-to-use probabilistic calibration of reduced-form IAM parameters. Mapping IPCC assessments to a stochastic DICE specification typically requires additional modeling choices (e.g., selecting a parametric family, matching assessed quantiles, and maintaining internal consistency of the reduced-form climate emulator when varying physical parameters). In this paper we therefore (i) retain the benchmark specification for our main numerical experiments, and (ii) explicitly compare key benchmark inputs to AR6 assessed ranges where a direct comparison is meaningful (see the discussion of ECS below and ipcc2021wgi_spm).
Finally, SSP/RCP scenario analysis is naturally viewed as a complementary “between-scenario” layer: one may recalibrate the exogenous trajectories (population, productivity, emissions intensity) to a given SSP pathway and then re-solve the resulting stochastic control problem with the same deep LSMC algorithm; see, e.g., riahi2017ssp and the ScenarioMIP/CMIP6 overview oneill2016scenariomip.
@@
Productivity growth. Assuming a Cobb-Douglas production function, the growth in total factor productivity $A$ models the growth in output that is not explained by growth in inputs of labor and capital used in production. The DICE model assumes $A$ evolves according to $A_{t+1}=A_{t}/(1-g_{A}(t))$, where $g_{A}(t)$ is the deterministic growth rate which is specified in Table 1. nordhaus2018 assumes $g_{A}(0)$ is normally distributed with mean $0.076$ and standard deviation $0.056$. But in this case, using the dynamics for the growth rate, we can model $g_{A}(t)$ as normally distributed with mean $g_{A}(0)\exp(-0.005t\Delta)$ and standard deviation $0.056\exp(-0.005t\Delta)$. In order to remove extreme cases, we truncate this distribution at the mean $\pm$ two standard deviations. The evolution of $A_{t}$ is shown in Figure 1.
@@
Abbildung 1: Evolution of total factor productivity AA under the assumption that the growth rate $g_{A}$ is uncertain.
@@
The rate of decarbonization. Uncontrolled industrial $\mathrm{CO}_{2}$ emissions are given by a level of carbon intensity, $\sigma_{t}$, times gross output. The DICE model assumes $\sigma$ evolves according to $\sigma_{t+1}=\sigma_{t}\exp(g_{\sigma}(t)\Delta)$, with a deterministic growth rate $g_{\sigma}(t)$ which is specified in Table 1. nordhaus2018 assumes $g_{\sigma}(0)$ is normally distributed with mean 0.0152-0.0152 and standard deviation $0.0032$. We therefore model $g_{\sigma}(t)$ as normally distributed with mean $g_{\sigma}(0)(1-0.001)^{t\Delta}$ and standard deviation $0.0032(1-0.001)^{t\Delta}$, truncating the distribution at the mean $\pm$ two standard deviations in order to remove extreme cases. The evolution of $\sigma_{t}$ is shown in Figure 2.
@@
Equilibrium climate sensitivity (ECS). Equilibrium climate sensitivity (ECS) measures the long-run increase in global mean surface temperature following a doubling of atmospheric $\mathrm{CO_{2}}$. The DICE-2016R2 benchmark corresponds to ECS $3.1\degree\mathrm{C}$ for an equilibrium $\mathrm{CO_{2}}$ doubling (see Table 1 and the denominator in the definition of $\psi_{2}$). nordhaus2018 models ECS as a log-normal distribution, $\exp(X)$ with $X\sim\mathcal{N}(1.1060,0.2646^{2})$, and we follow this benchmark specification, truncating at the mean $\pm$ two standard deviations.
[…]
Because ECS enters through the reduced-form DICE climate emulator, varying ECS should ideally be accompanied by a consistent re-calibration of the remaining climate-module parameters (e.g., feedback and heat-uptake coefficients); see folini2025climate for discussion. In our benchmark numerical study we follow nordhaus2018 by varying ECS through the DICE climate-sensitivity parametrization while keeping the remaining emulator coefficients at their DICE-2016R2 values; our UQ results should be interpreted conditional on this benchmark calibration.
The damage function. The DICE model assumes that climate-induced economic damages are a quadratic function of the increase in atmospheric temperature. It is modeled as a fractional loss of global output from greenhouse warming, $d_{t}=\pi_{2}\times(T_{t}^{\mathrm{AT}})^{2}$, where $\pi_{2}$ denotes a damage coefficient representing the severity of the economic impact of global warming. The DICE model assumes $\pi_{2}$ to be equal to 0.00236. nordhaus2018 models the $\pi_{2}$ by a normal distribution with mean 0.00236 and standard deviation 0.00118. We use the same distribution but truncate it at the mean minus one standard deviation, and at the mean plus two standard deviations.
The functional form and calibration of damages in IAMs are actively debated, and alternative specifications can imply materially different SCC levels and policy prescriptions; see, e.g., (botzen2012how; carleton2022valuing; dietz2024iam). Our contribution here is methodological: the deep LSMC recursion applies unchanged under alternative damage specifications and recalibrations.
The carbon cycle. The carbon cycle coefficient models the equilibrium concentration of carbon in the biosphere and upper level of the oceans. The DICE model assumes it to be equal to 360 gigatonnes of carbon (GtC). In Table 1, it corresponds to the value 360 appearing in the definitions of $\phi_{12}$ and $\phi_{23}$. nordhaus2018 models this coefficient as a log-normal distribution, $\exp(X)$ with $X\sim\mathcal{N}(5.8510,0.2649^{2})$. We do the same, truncating at the mean $\pm$ two standard deviations.
Abbildung 3: Density plots of the parameter distributions of equilibrium climate sensitivity (left panel), the damage coefficient (middle panel), and carbon cycle coefficient (right panel).
@@
  • Assuming $\pi_{2}\sim\mathcal{N}(0.00236,0.00118^{2})$ implies a roughly $2.3\%$ probability of $\pi_{2}$ being negative. This is a non-negligible scenario. Given that the DICE model aims to combine equations for the economy and climate, it is highly questionable to assume the damage coefficient could be below or just above zero. Moreover, the assumption of a log-normal distribution for the equilibrium climate sensitivity and the carbon cycle coefficient also entails a non-negligible probability of those coefficients being close to zero.
    nordhaus2018 avoids this issue by discretizing the distributions, separating them into quintiles, and then calculating the expected values of the random variables within those quintiles. These expected values are taken as realizations of discrete uncertain variables, yielding sufficiently positive lowest realizations for the coefficients. Inspired by this approach, we also truncate the distributions of the random variables, however, without discretizing them. All quantitative results below should be interpreted conditional on this benchmark truncation, which avoids issues with too low damage coefficients and temperature sensitivities, as well as extreme growth rates for total factor productivity and carbon intensity.
@@ 3 The deep least-squares Monte Carlo method
The numerical solution of the model is achieved using the endogenous state least-squares Monte Carlo (LSMC) algorithm with control randomization, as introduced by kharroubi2014numerical and adapted for expected utility optimal stochastic control problems by andreasson2022. This method approximates the conditional expectation of the value function in the Bellman equation using regression with a quadratic loss function applied to the transformed value function. Typically, regression basis functions are ordinary polynomials of the state and control variables, usually up to the third order. In our implementation, we use deep neural networks to approximate the regression predictor. To mitigate transformation bias in the regression estimate of the conditional expectation, we employ the smearing estimate (see Duan1983Smearing; andreasson2022).
Remark 3.1.
The transformation-bias adjustment and smearing-based correction used in our regression step (see Subsection 3.1) follow the approach of andreasson2022. Our methodological contribution here is the extension of this baseline to a stochastic climate-economy control problem with continuous, inequality-constrained policies and multiple simultaneous uncertainty sources: we use deep neural networks as the regression class, and we incorporate IAM-specific design choices (post-decision-state formulation, normalization, and low-discrepancy sampling) that materially improve numerical stability and precision in the stochastic DICE setting.
@@
where $Z_{0},Z_{1},\ldots,Z_{N-1}$ are independent disturbance terms, i.e. the state of the next period depends on the current state’s value, the current period’s control decision, and the realization of the disturbance term. This discrete-time representation also covers models specified originally in continuous time: after a time discretisation of a controlled SDE one obtains a one-step transition of the form (7), with $Z_{t}$ representing (for example) Brownian increments over a time step.
@@
$V_{t}(x)=\sup_{\pi_{t}\in\mathcal{A}_{t}}\bigg\{R_{t}(x,\pi_{t})+\mathbb{E}\big[\rho V_{t+1}(X_{t+1})\,\big|\,X_{t}=x;\pi_{t}\big]\bigg\},\quad t=N-1,N-2,\ldots,0,$ (8)
where the expectation is conditional on the state $X_{t}=x$ and the policy $\pi_{t}$ at time $t$. For further details on dynamic programming, we refer the interested reader to the excellent monograph by fleming06 on the subject.
@@
Here, $\mathcal{A}_{t}$ denotes a set of admissible values of $\pi_{t}$, which may depend on $x$. When the number of stochastic variables is more than three, it usually becomes computationally infeasible to use quadrature-based methods to evaluate the conditional expectation in (8), making simulation methods like LSMC preferable. The increase in the number of state variables will call for LSMC too.
@@
Remark 3.2.
The regression step to approximate (10) is closely related in spirit to regression-based expectation approximations used in computational economics, most notably the parameterized-expectations approach of den1990solving and its neural-network implementations (e.g. duffy2001approximating), as well as recent ML projection methods (e.g. valaitis2024machine). They parametrize conditional expectations that appear in equilibrium conditions and learn them via forward in time simulation with regression. A key difference of our approach is that we solve a recursive stochastic control problem backward in time by approximating the continuation value conditional expectation in the Bellman equation via regression and then recover the optimal control by a separate numerical maximization step. This separation is convenient in stochastic DICE because controls are constrained and the state dynamics are endogenous.
@@ 3.1 Transformation bias and heteroskedasticity
For $\alpha\neq 1$, the inverse is $H^{-1}(y)=\frac{1}{1-\alpha}\log((1-\alpha)y)$, which is well-defined on the domain $\{y\in\mathbb{R}:(1-\alpha)y>0\}$. In particular, for $\alpha>1$ one has $H(x)<0$ for all $x$, so $H^{-1}$ is applied to negative arguments throughout.
@@
Under the homoskedasticity assumption that the distribution of $\epsilon_{t}$ does not depend on $(X_{t},\pi_{t})$, the mean of the transformed residuals in (16) is constant in $(X_{t},\pi_{t})$ and can be precomputed and reused.
If heteroskedasticity is present in the regression with respect to the state and control variables, a method that accounts for heteroskedasticity is required. In this case, the conditional variance can be modeled as a function of covariates:
@@
where $\{\bm{g}_{\theta}\colon\theta\in\widehat{\Theta}_{t}\}$ is another parametrized family of predictor functions. There are various standard methods to estimate $\bm{g}_{\theta}$ and the smearing estimate with controlled heteroskedasticity can then be used as discussed in andreasson2022.
@@
  • The method presented in Algorithm 1 is called the regression surface approach. A common alternative is the realized value approach, where the value function $V_{t}(x)$ in Equation (8) is not computed by using the approximation of the conditional expectation (which was needed to find the optimal policy according to Equation (9)), but rather by computing the discounted sum of rewards along one trajectory starting from the state $x$ at time $t$. While promising greater numerical stability than the regression surface approach, the realized value approach requires calculating optimal decisions along the individual trajectories, which comes at a significant computational cost. For details on this approach, we refer to andreasson2022 and references therein. Originally, we also implemented the realized value approach, however, we found that the regression surface approach provided a sufficiently accurate solution for the number of sample points chosen in our numerical study in Section 4.
@@ 3.2 Neural networks
In our paper, we choose for the parametrized family of functions $\{\bm{f}_{\theta}\colon\theta\in\Theta\}$ the class of deep neural networks. This algorithmically generated class of functions has found tremendous success in all fields of science. Over the years, it has been shown that neural networks can act as surrogate functions in many models, due to their far-reaching approximation capabilities.
In classical LSMC implementations for Bermudan/American option pricing, the regression model for the continuation value is often taken to be a low-degree polynomial in the state variables (e.g. longstaff2001valuing; tsitsiklis2001regression; belomestny2010regression). This can be effective when the continuation value is sufficiently smooth and the effective dimension is small. We also note that in Bermudan option pricing, the state dynamics is not affected by control and thus continuation value does not depend on control helping to get good results with a simple polynomial approximation.
In the stochastic DICE setting, the continuation value displays strong nonlinearities (utility curvature, nonlinear damages, multiplicative growth/climate components) and the mitigation control is constrained to $[0,1]$, which can produce kinked policy rules and extended boundary regions. Polynomial bases may then require higher degrees and many interaction terms to achieve a robust fit; the number of monomials of total degree at most $p$ in $d$ variables equals $\binom{d+p}{p}$ and grows rapidly with $d$ and $p$, which can lead to ill-conditioned least-squares problems and sensitivity to basis selection. Deep neural networks provide a flexible nonparametric alternative that avoids committing to an ex ante functional form and have been used as function approximators in regression-based dynamic programming methods (tsitsiklis2001regression; valaitis2024machine). Another reason for simple polynomial approximation may not work well for estimation of mitigation control is because optimal control is found in the maximisation step after conditional expectation is approximated. If the functional form of polynomial approximation wrt control is not good, then this may lead to a very bad results for optimal control. ‘Non-parametric’ nature of neural network regression adapts to the correct shape of the function to approximate without guessing the functional form and thus should lead to better results.
Theorems that establish approximations are referred to as universal approximation theorems (UAT); notable contributions include cybenko89 and hornik91. These theorems establish the topological density of sets of neural networks in various topological spaces. One speaks of the universal approximation property (kratsios21) of a class of neural networks. Unfortunately, these theorems are usually non-constructive. To numerically find optimal neural networks, one typically combines backpropagation (see, for example, rumelhart86) with ideas from stochastic approximation (robbins51; kiefer52; dvoretzky56).
Assuming sufficient integrability, the conditional expectation in Equation (10) is the orthogonal projection of $\rho V_{t+1}(X_{t+1})$ onto the subspace spanned by $(X_{t},\pi_{t})$ in the space of square-integrable random variables. The universal approximation property of neural networks in this space (see, for instance, hornik91) then justifies the approximation of $\Phi_{t}(X_{t},\pi_{t})$ by $\bm{f}_{\theta}(X_{t},\pi_{t})$ for a suitably chosen neural network $\bm{f}_{\theta}$.
@@
[…]
@@
4:   $\widehat{V}_{t}(\widetilde{X}_{t})\coloneqq R_{t}(\widetilde{X}_{t})$